### v1

# Quantum-inspired Techniques in Tensor Networks for Industrial Contexts

Abstract
In this paper we present a study of the applicability and feasibility of quantum-inspired algorithms and techniques in tensor networks for industrial environments and contexts, with a compilation of the available literature and an analysis of the use cases that may be affected by such methods. In addition, we explore the limitations of such techniques in order to determine their potential scalability.
1 Introduction
Quantum computing is a novel field that is receiving considerable interest due to its particular way of performing operations and processing information, being able to tackle highly complex computational problems. In the information and digitalization era in which we find ourselves, this technology has a remarkable applicability for industrial cases, for example in the optimization of delivery routes [1, 2], packet storage [3] or quantum machine learning [4] applied to various contexts [5, 6, 7, 8].
However, despite the great progress in the development of digital quantum (or ‘logic gate’) computers, substantial challenges remain, such as the fragility of quantum states, errors in quantum operations, or the scalability of the number of qubits.
In this context, there is a need to explore alternative approaches that can take advantage of the computational properties of quantum systems, but do not have to run on digital quantum computers. One such approach may be quantum annealing [9, 10], which runs on specialized quantum devices dedicated to combinatorial optimization of particular cost functions. Another one is digital annealing [11, 12, 13], which draws on properties of quantum annealing to simulate it classically.
On the other hand, one of the most promising technologies, although more unknown within the industry, is tensor networks (TN) [14]. Tensor networks are a class of quantum-inspired algorithms and techniques based on mimicking the tensor operations performed by a quantum computer, but executing them on classical computers. By using tensor properties, the execution of such operations can be optimized, especially in cases where the entire quantum state vector is not required, but only properties of it.
Another relevant feature of tensor networks is that they allow the efficient representation of certain families of quantum states, by means of representations such as the matrix product state (MPS)[15, 16], also called tensor train (TT), or projected entangled pair states (PEPS) [16]. In this way, with a reduced amount of memory we can perform quantum calculations and obtain properties of complex systems. These same representations have been relevant to the world of machine learning, since they are capable of compressing models by reducing the memory required without losing a notable amount of precision [17, 18, 19].
Given all these capabilities, tensor networks are highly susceptible to be used in industrial contexts, being able to address highly complex and large problems efficiently. Despite being a relatively mature field at the academic level (initiated in 1971 by Penrose [20]), its study applied to industry is recent and always associated with quantum computing. However, given the great results achieved in recent years by tensor networks, outperforming quantum algorithms executable on current hardware [21, 22], many companies are beginning to take an interest in the study and implementation of this technology.
In this paper we survey the main uses of tensor networks for industrial use cases, compiling state-of-the-art articles and exposing the strengths and weaknesses of each technique. In addition, we classify the algorithms according to their specific use case. The field of tensor networks has an extensive and complex literature, so the aim of this article is to serve as an introductory guide to it.
2 Introduction to tensor networks
The first thing we will do is to make a brief introduction to tensor networks in order to understand the subsequent sections. In an effort to stay within the main objective of the paper, we will assume knowledge of basic linear algebra concepts and focus on tensor networks as such. Complementary information on tensor networks can be found in [14, 23].
A tensor network (TN) is a graphical representation of a multilinear equation between different tensors. We can see an example in Fig. 1, where we have the tensor network representing the 2-tensor $T_{ip}$ whose elements are obtained by means of the contraction operation
$T_{ip}=\sum_{j,k,l,m,n,o}A_{ijkl}B_{jm}C_{kmn}D_{lo}E_{nop}.$ (1)
Figure 1: Tensor network representing the 2-tensor $T_{ip}$ of Eq. 1.
As can be seen, this is a very simple representation of a very complex operation, in whose larger generalizations it is complicated to determine the relationships between the tensors contracted and the most efficient way to contract them. In addition, we can further abstract the tensor network by eliminating the names of the indexes and tensors if we know by some rule what they should be. An example would be that the tensor indexes are placed according to the clockwise orientation.
The rules followed by such representations are as follows:
  1. 1.
    Each node of the network represents a tensor.
  2. 2.
    Each outgoing line of a tensor represents an index of that tensor. Therefore, a node with $n$ outgoing lines represents an $n$-tensor.
  3. 3.
    Each line connecting two tensors represents that both tensors are contracted with respect to a common dummy index.
  4. 4.
    Lines that only come out of one tensor, but do not connect to another, represent free indexes. Therefore, a tensor network with $n$-free indexes represents an $n$-tensor.
  5. 5.
    Nodes that are not connected by an index are in tensor product with each other.
A highly interesting and useful type of tensor network is the already mentioned MPS [15], since it consists of an one-dimensional and decomposed representation of a tensor by means of the product of matrices. We can see an example in Fig. 2 a), where we see the MPS representation of a 5-tensor. This MPS representation is composed of 2 types of indexes: the bond indexes, which will have dimension $b$, and the physical indexes, of dimension $d$, corresponding to the dimensions of the tensor to be represented. The physical indexes are the free indexes, which coincide with the indexes of the tensor to be represented, while the bond indexes are those that connect the tensors of the network to each other and are responsible for accounting for the correlations between the physical indexes. It should be noted that there are other versions of the MPS representation with different boundary conditions (the way to link the end nodes), but the simplest is this one. There is also the possibility that the dimensions of the bond indexes are different, as are those of the physical indexes, for reasons we will mention at the end of this section.
For an $N$-tensor MPS representation, with constant bond dimension $b$ and constant physical dimension $d$, the number of elements to store is $2db+(N-2)db^{2}$ versus the $d^{N}$ that would be needed in the dense, contracted version of the tensor, $db$ for each end node and $db^{2}$ for each node in the interior of the chain.
Any tensor can be represented in MPS representation exactly, although not always efficiently, since $b$ can be of the order of $d^{N}$ in certain complex cases and where the correlation is very strong. Therefore, it will be important to consider how $b$ increases with the size of the tensor to be represented. Still, many cases of interest can be represented in MPS representation efficiently. For example, in academic contexts its most widespread use is that of representing quantum states of $N$ particles, that is, state vectors of $d^{N}$ components.
An extension of the MPS representation is the Matrix Product Operator (MPO) [24], consisting in allowing each node to have 2 physical indexes, as we see in Fig. 2 b). The usefulness of this representation lies in cases where we want to represent a matrix of $d^{2N}$ components, such as a quantum interaction. In this case, the number of elements to store is $2d^{2}b+(N-2)d^{2}b^{2}$.
Figure 2: a) MPS form of a 5-tensor, b) MPO form of a 10-tensor.
Such forms can be obtained mainly in three ways. The first is by logical construction of the representation, so that the interactions between the different nodes of the tensor network generate the tensor we want. The second will be the variational construction, consisting of starting with a test set of tensors and using techniques such as gradient descent to get it to become the one we want to represent. The third is by singular value decomposition (SVD) [25], iterative, where we perform an SVD for each pair of physical indexes. There is also an optimized method for sparse tensors [26].
Importantly, either representation can be used to represent a vector, matrix, or general tensor, since we can convert an $m$-tensor to an $n$-tensor by techniques such as ‘grouping’ and ‘splitting’ [14], a bijective mapping from one set of indexes to another set of indexes.
Consider an example of obtaining the MPO representation of $n$ nodes of a $N\times M$ matrix. To do this we will have to start by transforming the matrix into a $2n$-tensor, by means of a splitting. The resulting $A^{\prime}$ tensor will have dimensions $\vec{D}$ such that
$N=\prod_{i=0}^{n-1}D_{i},\qquad M=\prod_{i=n}^{2n-1}D_{i}.$ (2)
Now we join the pairs of input and output indexes that we want to be part of the same node of the MPO form and we group them to obtain a tensor $A^{\prime\prime}$ whose new dimensions $d_{i}=D_{2i}D_{2i+1}$ are $d_{i}=D_{2i}D_{2i+1}$. Now we perform the iterative process of searching for the MPS representation of that tensor, to finally split the physical indexes of each node to obtain the original input and output indexes. The whole process can be seen in Fig. 3.
Figure 3: Process of compression of a matrix to its MPO representation. 1) We apply a splitting. 2) We join the indexes of each node in pairs. 3) We apply a grouping of the index pairs. 4) We perform the iterative SVD. 5) We apply a splitting to each physical index.
If instead of wanting an exact representation of a tensor, we would like an approximate representation of it in order to have a smaller $b$ and thus save memory, we can take advantage of these two representations. In the variational construction we will simply operate with smaller $b$, while in the iterative SVD we will perform truncated SVD (TSVD) [27], keeping only the $b$ largest singular values, which represent the most significant correlations. We can also approximate a previously obtained MPS representation [28]. The accuracy of this approximation will depend both on the tensor itself that we want to represent and on the mapping we use, since if we put together in the same nodes the most correlated ‘columns’, all their correlations will be within the node itself. Moreover, if we have the most correlated indexes adjacent to each other, we will need a smaller bond dimension, since the intensity of the correlations between two nodes of an MPS representation decays exponentially with the geodesic distance between them, the minimum distance that separates them.
3 Basic tensor properties and more complex techniques
In this section we will make a connection between the basic properties of tensor networks and certain more advanced and complex techniques. In this way we will create a bridge between the abstraction of tensor networks and their practical applications. To do so, we will approach each case explaining which properties are determinant in it and which is the advantage it offers with respect to other methods.
3.1 Simulation of quantum systems
As is evident from their historical context, tensor networks are especially intended for the efficient representation of many-body physical systems. This is of great interest in fields such as condensed matter physics [29, 30] or quantum computing [31, 32].
For this purpose an estimated (or ‘ansatz’) solution that is representable as a tensor network is chosen and worked with, translating into tensor notation also physical operators and magnitudes. This allows us to work with a limited amount of memory, unlike the case of working with the entire statevector. Cases where tensor networks are very successful are in obtaining minimum energy states in quantum chemistry [33] and in the time evolution of certain quantum states to obtain certain magnitudes and properties of them [21, 22].
3.2 Machine learning model compression
The best known application of tensor networks is the compression of machine learning models [17, 18, 19]. This can be done either by compressing an already created model and training it or by creating a compressed model and training the tensor network that represents it directly. The key to model compression lies in finding an efficient representation of the model that preserves its analysis properties. We will explain how it works by compressing a dense layer in a neural network to its approximate MPS and MPO representations.
To begin with, we will have to keep in mind that the action of a dense layer of a neural network, prior to the action of an activation function, can be written as
$\vec{v}=A\vec{x}+\vec{c},$ (3)
where $A$ is a matrix of dimension $N\times M$, $\vec{v}$ and $\vec{c}$ are vectors of dimension $N$ and $\vec{x}$ is a vector of dimension $M$. The input to the layer is the vector $\vec{x}$, while the output will be the vector $\vec{v}$. The layer will be characterized by the $M$ components of the matrix $A$ and the $N$ components of the vector $\vec{c}$, which will be the ones to be trained. Therefore, the layer will have a number of $N(M+1)$ parameters. To reduce the number of parameters we can represent the matrix $A$ as an approximate MPO form with a certain bond dimension $b$ so that we have a smaller number of parameters than in the original dense version. We can do the same with the vector $\vec{c}$ obtaining its approximate MPS form. In this way, we have compressed the model layer. This is favorable both to save memory space and to train the model faster and avoid problems such as overfitting, depending on the representation used and the original model.
Since the $A$ matrix approximation and the $\vec{c}$ vector approximation can be made arbitrarily close to the originals by increasing the $b$ bond dimensions, we can always start from the original model and compress it until we saturate a preset tolerable error. However, we can also compress the elements of the model and run it to see how accurate it is with respect to the original model, increasing the compression until we reach that maximum error.
Another possibility would be to use a compressed version of the untrained model, with a controlled initialization as in the original, and train the compressed model directly. This has the advantage that we do not lose accuracy by approximating our model, but train an already pre-approximated model. We can also initialize the model by compressing an already trained model and retraining it.
There are many compression schemes in addition to the MPS and MPO, such as the PEPS [34, 35] or the MERA [36]. These allow to efficiently perform certain operations on their compressed representations and to apply various optimization methods, such as the Density Matrix Renormalization Group (DMRG) [36, 37].
3.3 Application of high dimensionality operations
A problem that we may have in real industrial cases is the need to apply high dimensional operations, such as creating certain kernels for methods like the Support Vector Machine [38]. We can take as an example wanting to apply a kernel in tensor product of the components of an $\vec{x}$ input vector of $N$ components [39]. This kernel is
$\phi(\vec{x})_{i_{1},i_{2},\dots,i_{N}}=\phi_{1}(x_{1})_{i_{1}}\phi_{2}(x_{2})_{i_{2}}\dots\phi_{N}(x_{N})_{i_{N}},$ (4)
where each kernel $\phi_{j}(\cdot)$ affects the $j$-th component of the vector $\vec{x}$. In this case, we can consider a kernel $j(\cdot)$.
$\phi_{j}(x_{j})=(x_{j},1),$ (5)
so that the global kernel $\phi(\cdot)$ will give us all possible products between the components of the input vector, as a $2^{N}$ component tensor to express correlations in the vector. As we can see, in the standard way this would require an exponential amount of memory and time in the size of the input vector. Furthermore, if we want to apply a matrix to such a kernel we would need a matrix $A$ of dimension $M\times 2^{N}$, which scales prohibitively.
However, by taking advantage of the properties of tensor networks we can use an MPO representation for such a matrix, of the form depicted in Fig. 4, where only $2N$ smaller tensors need to be contracted.
Figure 4: Eq. 4 and MPO layer for an input vector of $N=5$ components.
Figure 5: Contraction of the first three nodes of the MPO layer with the first three nodes of the kernel. 1) We contract a kernel node with its corresponding MPO layer node. 2) We contract the next kernel node with its layer node. 3) We contract the two new nodes obtained. 4) We repeat the contraction process for the next pair of nodes. 5) We contract the current nodes, maintaining the physical index.
With this methodology we can apply an operation in which we weight elements of type $x_{1}x_{2}x_{3}x_{4}x_{5}$, $x_{1}x_{2}x_{3}x_{4}$, $x_{1}x_{2}x_{3}$, $x_{1}x_{2}x_{2}$ and $x_{1}$, efficiently and with a reduced number of elements.
3.4 Data compression and denoising
In the field of big data there is a need to work with large volumes of data, which can be very expensive or unfeasible for certain systems. The tensor compression property that we have seen in representations such as MPS allow us to compress [40] and handle [41] large amounts of data. To do this we just need to convert the data into a tensor and compress it. In addition, if we allow ourselves to reduce the precision of its elements or are only interested in its global characteristics, we can use approximate representations of the tensor to save memory space.
3.5 Combinatorial optimization
Solving most combinatorial optimization problems requires runtimes exponentially increasing with problem size. Given the properties of quantum computation, many can be tackled through it efficiently. A similar approach to quantum lies in the use of tensor networks for combinatorial optimization. There are different types of combinatorial optimization solvers.
The first one is the one developed in [42], running in a black box. This algorithm starts from a cost tensor, $T$, whose subscripts are the components of the solution vector. This tensor will have the cost of all possible states, although it is not necessary to know the tensor explicitly: it is enough to have a function that gives us the cost as a function of its indexes. With this tensor, we will make an approximation of its max volume submatrix, the submatrix containing its largest singular values in product, by means of an MPS representation. From this representation we will obtain, with high probability, the lowest cost element.
The problem with this algorithm is that when we work with problems with many invalid configurations, it tends to fail because it sees that almost all or all the elements are the same, since it is sampling the tensor. Therefore, we need an encoding, however complicated it may be, in which we reduce the space of invalid configurations, so that the algorithm sees a certain functional map.
The second possibility is based on the use of Generative Artificial Intelligence (AI) boosted with tensor networks [43].
The third possibility is the use of imaginary time evolution to a quantum state in uniform superposition of all possible combinations, equivalent to the application of a Hadamard gate to each qubit of a quantum circuit. If from this uniform superposition we damp each combination exponentially to its associated cost, the optimal combination will be the one with the largest amplitude, resulting in a higher measurement probability.
Due to its non-unitareity, at the quantum level this type of operation cannot be performed, but at the classical level it is possible to perform [44]. If we needed to take into account all possible combinations, we would need an unbearable amount of memory, resulting in a brute-force search. However, by combining the ability to compress information and simulate quantum states with the ability to apply non-unitary tensor network operations, we can create such a state for various problems [45, 46, 47]. This can be done in an exact or approximate way, by recompressions, depending on how much memory we want to devote to it.
The main limitation is that for NP-Hard problems, the most complex ones, we will need an exponential amount of memory and time to solve them exactly. By making intelligent adjustments in the encoding and expression of the problem, the creation and contraction of the tensor network and applying various simplifications in the constraints, we can create efficient approximate algorithms from the exact ones, avoiding exponential scaling.
The fourth is the simulation of a physical system represented by a variational tensor network, with a cost function given as a Hamiltonian that we can write in MPS or MPO representation. In this case we will perform variational updates of the state tensor network, using techniques such as gradient descent or DMRG to minimize the energy of the system to find the optimal combination. The effectiveness of these types of techniques depends very much on the problem to which they are applied and they can get stuck in local minima, especially for problems with constraints. In addition, there is the limitation that we must be able to express the cost function in a tensor network representation.
4 Use case studies
In this last section we will present the main use cases in which we can see a direct and significant potential of tensor networks techniques to industrial and business contexts.
4.1 Finance
In the contemporary financial domain, high dimensionality emerges as a challenging phenomenon. With rapid technological advancement and the availability of large-scale data, financial markets have become inherently multidimensional. Thanks to this, in-depth market analysis is possible.
4.1.1 Portfolio optimization
Portfolio optimization consists of determining the best allocation of financial assets in which to invest in order to maximize a cost function, which is usually one that maximizes return and minimizes risk. This problem is already complicated in a static way, becoming extremely complex in a dynamic way, with a portfolio to be optimized over a series of trading days. The high dimensionality and constraints of the problem make it generally an intractable problem in an exact manner.
Tensor networks can be used to explore the space of inversion solutions in order to search for the optimal combination. In [48] we can see how they use imaginary time evolution on an MPS representation to solve the optimization problem. They achieve sharpe ratio results, a measure of excess return per unit of risk on an investment, superior to those obtained by other methods for large sizes, although with lower payoff and much longer run time for large problems than Dwave’s hybrid solver.
4.1.2 Interpretable predictions
In [49], a recurrent neural network method is presented for making interpretable predictions, those in which the underlying reasoning can be understood, using an MPS representation of the recurrent layer weight matrix. It can be observed that the accuracy, Sharpe and return results are higher in the tensor network version than in the other compared versions.
4.2 Medicine
The field of medicine is one of the fields with the highest dimensionality, both for the diagnosis and treatment of diseases, as well as in the design and production of drugs. Because of this, there has recently been a great interest in the application of Artificial Intelligence techniques in this field, obtaining great results.
4.2.1 Drug discovery
Drug simulation is a monumental challenge in contemporary science due to the complexity of molecular interactions and the enormous number of possible configurations. In this context, quantum computing offers a clear advantage as it is able to model and treat phenomena such as the electronic structure of molecules and interactions between atoms more adequately than classical computing. As explained above, the ability of tensor networks to efficiently represent quantum states may allow us to simulate certain molecules that are currently not possible with quantum computers.
Similarly, we can save time and costs of new drug discovery by predicting new therapeutic properties associated with new drugs. This can be achieved in several ways, one of them being the tensor decomposition [50]. To this end, we create a drug-gene-disease tensor, which accounts for the relationships found between these drug-gene, drug-disease and gene-disease pairs. With this, we formulate a tensor completion problem. The initial tensor will have only the direct relationships of the pairs we have found, while the rest of the more complex relationships will be the ones we have to determine. For this, we will perform a generalized tensor decomposition (GTD), to model part of these relationships, and we will apply a multilayer perceptron (MLP) to capture the rest of the relationships.
The main limitation of this method is that its predictive capacity is inferior for new diseases, so other techniques must be coupled to it.
4.2.2 Medical image analysis
The analysis of images, and especially medical images, is a very complicated task, either because of the characteristics of these images, the noise in them or the high dimensionality that the problem can reach. However, as we have seen, tensor networks are ideal for cases with high dimensionality and noise.
In [51] authors present methods for classifying medical images, using exactly the scheme we explained in Ssec. 3.3, interpreting the images as an input vector and a feature map cosenoidal, and the MPO layer as a classifier. Thus, by contracting the tensor network we will obtain a vector that will tell us the probabilities of that image belonging to one group or another. This is done in combination with other more advanced machine learning techniques and finally we obtain an extremely high level of compression with levels of area under the ROC (Receiver Operating Characteristic) curve similar or superior to the methods with which it is compared. All this with much lower GPU requirements.
Medical image segmentation is vital for disease diagnosis, treatment monitoring, surgical guidance, experimental drug development and medical education. A method for medical image segmentation with tensorized transforms, introduced in the paper, is shown in [52]. In this method, a tensorized version of the self-attention module, a step of the transformer, is realized.
4.3 Simulation of quantum materials and topological materials
Simulation of quantum materials and topological materials is an extremely complex problem, because they are high-dimensional problems. Quantum many-body systems require an exponential memory in the number of elements to represent their wave functions. The utility of these materials is enormous, given their curious applications, such as high-speed electronics, solar cells, quantum sensors, telecommunications and quantum computing.
To deal with this, concrete ansatz are proposed [53, 54, 55, 56] which accurately and efficiently represent the minimum energy states of these systems. We start with an initial state and perform various transformations, such as DMRG or imaginary time evolution, to bring it down to its minimum energy state. Once this state is obtained, we can easily obtain properties of the state by applying the tensors of the quantities to be measured.
4.4 Battery simulation
Battery simulation is a very useful problem, but it also requires the study of quantum many-body systems. For this we can also use ansatz in tensor networks [57, 58], with which we can obtain properties of the minimum energy states.
4.5 Optimization
Optimization problems in industrial environments are very important, as they may require the optimization of the order in which processes are performed, such as the assignment of times and tasks to machines in the Job Shop Scheduling Problem (JSSP). However, due to the problem size in these cases, it is very expensive to solve such instances due to the high number of combinations and the complexity of the cost functions.
4.5.1 Route Optimization
The Traveling Salesman Problem (TSP) is a widely studied historical problem, in which a route must be chosen that traverses all nodes of a certain network only once and offers the lowest cost associated with it, given by the costs of going from one node to another. Generalizations and particular cases of this problem model sharing problems, which are very important in industry.
For this purpose, [45] propose a TN method that can solve generalized TSP cases, with different variants. This method takes as variable the node in which one is at each step of the route and thus model the cost function as a Quadratic Unconstrained Discrete Optimization (QUDO) to the nearest neighbor with non-repetition constraints. Its basic mechanics resides in the imaginary time evolution described in [46] and the connection of a set of layers that impose the constraints, eliminating non-compatible combinations in the represented state. By modifying such constraint layers and the QUDO modeling, different generalizations can be addressed, such as being able to repeat up to $N$ times a certain node, that the start and end can be different or that we can only pass through each group of nodes once. All these cases are studied in [45].
The main limitation of this method is the exponential scaling it suffers with the number of nodes due to the constraints, but iterative methods such as the one explored in [47] are able to limit this scaling successfully.
4.5.2 Post assignment optimization
A particular case of generalized TSP is the assignment of jobs to workers. Assume a set of workers who are already assigned to a set of jobs and we have a set of vacancies. The objective will be to redistribute workers to the set of jobs, including filled and vacant jobs, such that we maximize a certain cost function and there are only 0 or 1 workers per job. This problem has been studied and treated at the end of [45], using the formulation of [59], which seeks to maximize a combination between the productivity of the job and the compatibility of the worker assigned to it.
4.5.3 Manufacturing Sequence Optimization
The optimization of manufacturing sequences is a very useful industrial problem. We have a set of products to be manufactured sequentially, such that there is additional manufacturing time when moving from manufacturing one type of product to manufacturing another. This time can be a time to change parts of a machine, configuration or materials. Therefore, we want to obtain the sequence that minimizes this overall additional time.
This case is studied in [45], since it can be seen as a TSP in which we can be at each node as many times as there are outstanding products of the class associated to that node in the initial set. With this perspective, a node represents a class rather than a product to optimize the constraint layers.
4.5.4 Machine task assignment optimization
Optimal assignment of tasks to machines can be modeled as the choice of task execution on a set of machines such that there is a constraint between the tasks performed by one and another. An example of a constraint would be the following: ‘if machine 1 does a cutting, machine 2 cannot do a peeling’. These constraints can be extracted from a historical record or from the business knowledge of the usual machine manager. In addition, each task will have its own execution time on its corresponding machine.
This problem is addressed in [47], where a TN method with imaginary time evolution and constraint layers like those of [45] is combined with a genetic algorithm and an iterative method to reduce the time and memory scaling of such algorithm, taking advantage of properties of this type of problems. The main limitation of the method is that, although it has a high success rate, for certain cases it cannot obtain satisfactory solutions for a given memory limit.
4.6 Big Data
As its name indicates, in the field of big data large volumes of data are handled, so we need specific systems to deal with them.
The work [40] presents different methods to deal with big data problems using tensor network decompositions, as well as other topics. On the other hand, in [41] a tensor network method for secret sharing in big data is presented, based on the decomposition into an MPS or a Tucker decomposition [60] of the original data tensor by means of consecutive TSVDs. Between each TSVD a tensor perturbation will be performed, by multiplying the new node obtained from the SVD by a $\Delta$ matrix, and the tensor to be decomposed in the next step by $\Delta^{-1}$. This achieves that the global tensor is the same, but the nodes of the representation are different. Finally, the first node of the representation is randomized. In this way, parts of the dataset can be stored in a distributed way while maintaining the security of the data, since no one will be able to obtain any information from the data until all the nodes of the representation are available. At the same time, it presents a way to perform distributed computing for big data.
4.7 Classification
As we have seen in the case of medical image classification, using methods such as the one presented, different types of images or general data can be classified. In addition, there are more general methods. On the one hand, the inspiration of quantum circuits in tensor networks methods to make classification circuits [61]. In this case, hierarchical trees are addressed.
On the other hand, we have the use of PEPS for image classification, where results similar to those of GoogLeNet, VGG-16 and AlexNet [62] are achieved by means of simple coupled tensor network systems.
4.8 Artificial Intelligence
As we saw in Ssec. 3.2, various models dedicated to artificial intelligence can be compressed using tensor network representations. Combining that with what was shown in Ssec. 3.3, we can deal with much larger and more complex problems. In addition, we can add various quantum machine learning techniques [4] to design new models with the features we need.
On the other hand, creating more compact models that can be inserted and trained directly on more modest devices, such as mobile devices, can allow experimentation with more models simultaneously and new applications.
We can also realize privacy-preserving machine learning techniques using tensor networks [63] techniques. This is of vital importance when working with sensitive data, such as medical records.
4.9 Cybersecurity
Cybersecurity is a very complex field, in which extraordinary measures are needed to protect users’ privacy and detect possible attacks [64]. Here we can encompass the techniques already seen above applied to this context.
4.10 Anomaly detection
Anomaly detection is a problem consisting of identifying unusual cases in a data set. This problem is very useful in detecting computer attacks. Because anomalous data can present many structures, having a large space of possibilities, it is required to model in a very effective way the normal data so that they can be identified with respect to the anomalous ones. This usually requires analyzing high dimensional data or applying very complex operations to it.
The work [39] discuss a supervised method precisely to be able to deal with this problem. By applying an ansatz product like the one we explained in Ssec. 3.3, and an operation in variational MPO that we will train, we can project the normal data to a hypersphere. In this way, the normal points will remain on the surface of the sphere, while the anomalous ones will go to the center or to the outside of the sphere. Moreover, the process can be performed very efficiently due to the structure of the tensor network to be contracted, scaling remarkably well with the size of the problem.
It has been tested with different datasets against different algorithms, giving good results. Its main limitation is that it is a supervised algorithm, in which we require a well-performed labeling prior for it to work properly. Even so, it can be extended to be part of an unsupervised algorithm by adding automatic pre-labeling.
5 Conclusions
We have seen that tensor networks are a set of techniques with great applicability to the industrial context in an immediate way, with proven advantages and an extensive literature on the subject. We have explained the main features that make them especially attractive for high dimensionality problems and machine learning. We have also highlighted several applied cases and some of the techniques that have been used in the state of the art to address them, as well as their advantages and disadvantages.
Acknowledgement
The research leading to this paper has received funding from the Q4Real project (Quantum Computing for Real Industries), HAZITEK 2022, no. ZE-2022/00033.